# Public-Key Quantum Money and Fast Real Transforms

### v1

@@ 4 A New Algorithm for $\mathsf{QHT}$
$\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{4\pi ty}{N}\Big{)}\lvert 0\rangle\lvert y\rangle\lvert b\rangle+\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{4\pi ty}{N}\Big{)}\lvert 1\rangle\lvert-y\rangle\lvert b\rangle.$
@@
$\lvert\phi_{1}\rangle$ $=\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\Big{(}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\lvert 0\rangle+\mathrm{cas}\Big{(}-\frac{2\pi a(N/2-y)}{N}\Big{)}\lvert 1\rangle\Big{)}\lvert y\rangle\lvert b\rangle$
$=\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\Big{(}\lvert 0\rangle+(-1)^{b}\lvert 1\rangle\Big{)}\lvert y\rangle\lvert b\rangle,$
where we used the fact that $\mathrm{cas}(-\pi a+2\pi ay/N)=(-1)^{a}\mathrm{cas}(2\pi ay/N)$ and $(-1)^{a}=(-1)^{b}$. Applying the Hadamard transform to the first qubit gives
@@
Next, we uncompute the first qubit using a cnot gate with the last qubit, then apply a Hadamard transform and a swap to obtain:
@@
The last sum matches exactly with (12), which represents the quantum Hartley transform of $\lvert a\rangle$. Therefore, we have successfully performed the operation:
@@
  9. 9.
@@
Algorithm 1 is correct and can be implemented using $\approx\log^{2}N+O(\log N)$ elementary gates.
@@
The correctness of the algorithm follows from the preceding discussion. Except for the unitary $U_{R}$ and the negation unitary of Step 4, all the steps in the algorithm can be implemented using $O(1)$ elementary gates. The negation in Step 4 can be implemented using $\approx\lceil\log N\rceil$ elementary gates. To implement the unitary $U_{R}$, which involves constructing the conditional operator $R(y,b)$ for arbitrary $y$ and $b$, we utilize the two-qubit operators
@@
Let $T(N)$ be the gate complexity of Algorithm 1 for an input state of dimension $N$. Assuming access to the gates $R_{j}$ as elementary gates, it follows from above that $T(N)\approx T(N/2)+2\log N+O(1)$. Therefore, $T(N)\approx\log^{2}N+O(\log N)$. ∎
@@ 4.1 Comparison With Other Algorithms
The recursive algorithm of [3] relies on the decomposition given in (8). We refer the reader to [3] for the definitions of $Q_{N}$ and $BC_{N/2}$, as well as their gate complexities. The permutation $Q_{N}$ can be implemented using approximately $\lceil\log N\rceil$ elementary gates. The implementation of $BC_{N/2}$ involves computing the two’s complement of an $(n-2)$-bit integer, along with applying the unitary $R(y,1)$ defined in (13). Each of these also requires approximately $\lceil\log N\rceil$ gates. Let $T(N)$ denote the gate complexity of the algorithm for input size $N$. Then Then $T(N)\approx T(N/2)+3\lceil\log N\rceil$, which implies that
$T(N)\approx\frac{3}{2}\log^{2}N+O(\log N).$
@@
It follows from the above and Theorem 4.1 that our algorithm for $\mathsf{QHT}_{N}$ has a gate complexity that is approximately 1.5×1.5$\times$1.5 × lower than that of [3], and approximately 2.5×2.5$\times$2.5 × lower than that of [22].
@@ 8 Computing The Serial Number Using Quantum Walks
$\lvert\phi_{j}\rangle=\frac{1}{\sqrt{\deg(j)}}\sum_{k=1}^{N}\lvert j,k\rangle,$

### v2

@@ 4 A New Algorithm for $\mathsf{QHT}$
$\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{4\pi ty}{N}\Big{)}\lvert 0\rangle\lvert y\rangle\lvert b\rangle+\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{4\pi ty}{N}\Big{)}\lvert 1\rangle\lvert N/2-y\rangle\lvert b\rangle.$
@@
$\lvert\phi_{1}\rangle$ $=\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\lvert 0\rangle\lvert y\rangle\lvert b\rangle+\frac{1}{\sqrt{N}}\Big{(}\lvert 1\rangle\lvert 0\rangle+\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}-\frac{2\pi a(N/2-y)}{N}\Big{)}\lvert 1\rangle\lvert y\rangle\Big{)}\lvert b\rangle$
$=\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\lvert 0\rangle\lvert y\rangle\lvert b\rangle+\frac{1}{\sqrt{N}}\Big{(}\lvert 1\rangle\lvert 0\rangle+(-1)^{b}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\lvert 1\rangle\lvert y\rangle\Big{)}\lvert b\rangle,$
where we used the fact that $\mathrm{cas}(-\pi a+2\pi ay/N)=(-1)^{a}\mathrm{cas}(2\pi ay/N)$ and $(-1)^{a}=(-1)^{b}$. Consider a controlled-cnot unitary $C_{X}$ that takes $\lvert c\rangle\lvert y\rangle\lvert b\rangle$ to $\lvert c\rangle\lvert y\rangle\lvert 1\oplus b\rangle$ if $c=1$ and $y=0$, and acts as identity otherwise. Then applying the unitary $HC_{X}H$ to the above state results in the state
$\frac{1}{\sqrt{N}}\sum_{y=0}^{N/2-1}\mathrm{cas}\Big{(}\frac{2\pi ay}{N}\Big{)}\Big{(}\lvert 0\rangle+(-1)^{b}\lvert 1\rangle\Big{)}\lvert y\rangle\lvert b\rangle.$
Applying the Hadamard transform to the first qubit gives
@@
Next, we uncompute the first qubit using a cnot gate with the last qubit, then apply a Hadamard transform to obtain:
@@
The final sum matches (12), except that the last qubit appears at the end rather than immediately after the ancilla qubit. This discrepancy can be resolved by relabeling the qubits throughout the recursion and applying a single permutation at the end of the computation. Therefore, we have successfully performed the operation:
@@
    Apply the controlled negation of Step 4
  7. 7.
    Apply the unitary $HC_{X}H$
  8. 8.
@@
  11. 11.
    Relabel the qubits to implement the effect of the swap $\lvert 0\rangle\lvert y\rangle\lvert b\rangle\mapsto\lvert 0\rangle\lvert b\rangle\lvert y\rangle$
  12. 12.
@@
Algorithm 1 is correct and can be implemented using $\approx 2\log^{2}N+O(\log N)$ elementary gates.
@@
The correctness of the algorithm follows from the preceding discussion. Except for the unitary $U_{R}$, the negation unitary of Step 4, and the multi-controlled cnot of Step 7, all the steps in the algorithm can be implemented using $O(1)$ elementary gates. The negation in Step 4 and the cnot in Step 7 can be implemented using $\approx\lceil\log N\rceil$ elementary gates. To implement the unitary $U_{R}$, which involves constructing the conditional operator $R(y,b)$ for arbitrary $y$ and $b$, we utilize the two-qubit operators
@@
Let $T(N)$ be the gate complexity of Algorithm 1 for an input state of dimension $N$. Assuming access to the gates $R_{j}$ as elementary gates, it follows from above that $T(N)\approx T(N/2)+4\log N+O(1)$. Therefore, $T(N)\approx 2\log^{2}N+O(\log N)$. ∎
@@ 4.1 Comparison With Other Algorithms
The recursive algorithm of [3] relies on the decomposition given in (8). We refer the reader to [3] for the definitions of $Q_{N}$ and $BC_{N/2}$, as well as their gate complexities. The permutation $Q_{N}$ can be implemented using approximately $\lceil\log N\rceil$ elementary gates. The implementation of $BC_{N/2}$ involves the following operations:
  1. 1.
    Compute a controlled two’s complement of a $(n-2)$-bit integer
  2. 2.
[…]
  3. 3.
    Apply the unitary $R(y,1)$ defined in (13).
  4. 4.
    Compute a controlled two’s complement of a $(n-2)$-bit integer
Each of the above steps require approximately $\lceil\log N\rceil$ elementary gates to implement. Let $T(N)$ denote the gate complexity of the algorithm for input size $N$. Then $T(N)\approx T(N/2)+5\log N+O(1)$, which implies that
$T(N)\approx\frac{5}{2}\log^{2}N+O(\log N).$
@@
It follows from the above and Theorem 4.1 that our algorithm for $\mathsf{QHT}_{N}$ has a gate complexity that is approximately 1.25×1.25$\times$1.25 × lower than that of [3] and [22].
@@ 8 Computing The Serial Number Using Quantum Walks
$\lvert\phi_{j}\rangle=\frac{1}{\sqrt{\deg(j)}}\sum_{k=1}^{N}\sqrt{P_{jk}}\lvert j,k\rangle,$